# Unveiling m-Sharpness Through the Structure of Stochastic Gradient Noise

### v1

@@ 6 Conclusion
In this work, we conducted a comprehensive theoretical analysis of SAM and its variants through an enhanced SDE modeling framework. Our findings reveal that the structure of SGN plays a crucial role in implicit regularization, significantly influencing generalization performance. Based on our analysis, we proposed Reweighted-SAM, an adaptive weighting mechanism for perturbation, which we empirically validated through extensive experiments. Our study provides a deeper understanding of the dynamics of SAM-based algorithms and offers new perspectives on improving their generalization performance.
@@
$b(x,\rho)=\sum_{m=0}^{\beta}\rho^{m}\,b_{m}(x)\;+\;O(\eta\rho^{\beta+1}),\qquad\sigma(x)=\sigma_{0}(x).$
@@
$L=\sum_{m=0}^{\beta}\rho^{m}\,A_{m}\;+A_{\Delta}\;+\;O(\eta\rho^{\beta+1})\,$,
@@ C.1 Mini-batch SAM
$\Sigma_{0,1}(X_{t}):=\mathbb{E}\bigl[(\nabla f_{\gamma}(X_{t})-\nabla f(X_{t}))\cdot\bigl(\frac{\nabla^{2}f_{\gamma}(X_{t})\nabla f_{\gamma}(X_{t})}{\|f_{\gamma}(X_{t})\|}-\mathbb{E}[\frac{\nabla^{2}f_{\gamma}(X_{t})\nabla f_{\gamma}(X_{t})}{\|f_{\gamma}(X_{t})\|}]\bigr)^{\top}\bigr].$
@@
    $\nabla f^{\mathrm{SAM}}\;=\;\nabla\Bigl(f+\rho\|\nabla f\|\Bigr)\quad\text{and}\quad\sqrt{\Sigma^{\mathrm{SAM}}}$
@@
Then, by Lemmas A.4, C.2, and D.2, all the conditions of Theorem A.2 are satisfied, and the proof is complete. ∎
@@
    $\nabla f^{\mathrm{m\text{-}SAM}}\;=-\nabla\bigl(f+\frac{\rho}{m}\mathbb{E}\|\sum_{\begin{subarray}{c}i\in\mathcal{I},\\ |\mathcal{I}|=m\end{subarray}}\nabla f_{i}\|\bigr)\quad\text{and}\quad\sqrt{\Sigma^{\mathrm{m\text{-}SAM}}}$
@@
Then, by Lemmas A.4, C.5, and D.2, all the conditions of Theorem A.2 are satisfied, and the proof is complete. ∎
@@ Appendix G Derivation of the Finite Difference Estimator in Equation (21)
$\mathbb{E}_{z}\!\Bigl[\Bigl(\tfrac{f_{i}(x+\delta z)-f_{i}(x)}{\delta}\Bigr)^{2}\Bigr]\;=\;\mathbb{E}_{z}\bigl[\bigl(\nabla f_{i}(x)^{\top}z\bigr)^{2}\bigr]\;+\;O(\delta)\;=\;\|\nabla f_{i}(x)\|^{2}+O(\delta).$
Thus the mean-squared estimate of the finite difference quotient differs from $\|\nabla f_{i}(x)\|^{2}$ by an $O(\delta)$ bias term. We compare two $d$-dimensional random vectors $z\in\mathbb{R}^{d}$:
@@ Appendix H Derivation of the Gibbs Distribution (20)
Appendix I Algorithm: Reweighted-SAM
Algorithm 1 Reweighted-SAM
@@ Appendix I Algorithm: Reweighted-SAM
ResNet-18 ResNet-50 WideResNet-26-10

### v2

@@ 6 Conclusion
In this work, we conducted a comprehensive theoretical analysis of SAM and its variants through an enhanced SDE modeling framework. Our findings reveal that the structure of SGN plays a crucial role in implicit regularization, significantly influencing generalization performance. Based on our analysis, we proposed Reweighted SAM, an adaptive weighting mechanism for perturbation, which we empirically validated through extensive experiments. Our study provides a deeper understanding of the dynamics of SAM-based algorithms and offers new perspectives on improving their generalization performance.
Acknowledgement
Trung Le, Mehrtash Harandi, and Dinh Phung were supported by the ARC Discovery Project grants DP230101176 and DP250100262. Trung Le and Mehrtash Harandi were also supported by the Air Force Office of Scientific Research under award number FA9550-23-S-0001.
@@ Acknowledgement
$b(x,\rho)=\sum_{m=0}^{\beta}\rho^{m}\,b_{m}(x)\;+\;O(\rho^{\beta+1}),\qquad\sigma(x)=\sigma_{0}(x).$
@@
$L=\sum_{m=0}^{\beta}\rho^{m}\,A_{m}\;+A_{\Delta}\;+\;O(\rho^{\beta+1})\,$,
@@ C.1 Mini-batch SAM
$\Sigma_{0,1}(X_{t}):=\mathbb{E}\bigl[(\nabla f_{\gamma}(X_{t})-\nabla f(X_{t}))\cdot\bigl(\frac{\nabla^{2}f_{\gamma}(X_{t})\nabla f_{\gamma}(X_{t})}{\|\nabla f_{\gamma}(X_{t})\|}-\mathbb{E}[\frac{\nabla^{2}f_{\gamma}(X_{t})\nabla f_{\gamma}(X_{t})}{\|\nabla f_{\gamma}(X_{t})\|}]\bigr)^{\top}\bigr].$
Remark C.1 (On normalization at critical points).
Note that SAM is ill-defined when the gradient is zero. To solve this, we may replace the denominator by $\|\cdot\|_{\varepsilon}=\sqrt{\|\cdot\|^{2}+\varepsilon^{2}}$ with a fixed $\varepsilon>0$, which is also a common implementation in practice. Under the standing assumption that $\nabla f$ is $L$-Lipschitz, the resulting coefficients are globally $O(L/\varepsilon)$-Lipschitz and $C^{1}$. All local Taylor or weak-approximation arguments and moment bounds in Appendix C continue to hold with constants depending on $\varepsilon$ but independent of $\eta,\rho$. The stated orders in $\eta,\rho$ are unaffected.
@@
    $\nabla f^{\mathrm{SAM}}\;=\;\nabla\Bigl(f+\rho\mathbb{E}\|\nabla f_{\gamma}\|)\quad\text{and}\quad\sqrt{\Sigma^{\mathrm{SAM}}}$
@@
Then, by Lemmas A.4, C.3, and D.2, all the conditions of Theorem A.2 are satisfied, and the proof is complete. ∎
@@
    $\nabla f^{\mathrm{m\text{-}SAM}}\;=\nabla\bigl(f+\frac{\rho}{m}\mathbb{E}\|\sum_{\begin{subarray}{c}i\in\mathcal{I},\\ |\mathcal{I}|=m\end{subarray}}\nabla f_{i}\|\bigr)\quad\text{and}\quad\sqrt{\Sigma^{\mathrm{m\text{-}SAM}}}$
@@
Then, by Lemmas A.4, C.6, and D.2, all the conditions of Theorem A.2 are satisfied, and the proof is complete. ∎
@@ Appendix G Derivation of the Finite Difference Estimator in Equation (21)
$\mathbb{E}_{z}\!\Bigl[\Bigl(\tfrac{f_{i}(x+\delta z)-f_{i}(x)}{\delta}\Bigr)^{2}\Bigr]\;=\;\mathbb{E}_{z}\bigl[\bigl(\nabla f_{i}(x)^{\top}z\bigr)^{2}\bigr]\;+\;O(\delta^{2})\;=\;\|\nabla f_{i}(x)\|^{2}+O(\delta^{2}).$
Thus the mean-squared estimate of the finite difference quotient differs from $\|\nabla f_{i}(x)\|^{2}$ by an $O(\delta^{2})$ bias term, implying that the estimator is approximately unbiased as $\delta\to 0$. We compare two $d$-dimensional random vectors $z\in\mathbb{R}^{d}$:
@@ Appendix H Derivation of the Gibbs Distribution (20)
Appendix I Algorithm: Reweighted SAM
Algorithm 1 Reweighted SAM
@@ Appendix I Algorithm: Reweighted SAM
ResNet-18 ResNet-50 WideResNet-28-10
@@
We present the experimental results of applying the proposed reweighting strategy to ASAM (Kwon et al.,, 2021) in Table 9. The results demonstrate consistent improvements, and further investigation into the effectiveness of applying the reweighting strategy to different SAM variants remains an interesting direction for future work.
Table 9: Test accuracy (%) comparison between ASAM and RW-ASAM on CIFAR-10 and CIFAR-100.
Method CIFAR-10 CIFAR-100
ResNet-18 ResNet-50 ResNet-18 ResNet-50
ASAM 95.86 $\pm$ 0.14 96.12 $\pm$ 0.23 79.17 $\pm$ 0.14 80.27 $\pm$ 0.33
RW-ASAM 96.02 $\pm$ 0.08 96.43 $\pm$ 0.17 79.46 $\pm$ 0.25 80.65 $\pm$ 0.16
Table 10: Hyperparameter settings for fine-tuning DistilBERT on GLUE tasks.
Task CoLA MNLI MRPC QNLI QQP RTE SST-2 STS-B WNLI
Batch size 32 64 32 64 64 32 64 32 32
Learning rate 2e-5 3e-5 2e-5 3e-5 3e-5 2e-5 3e-5 2e-5 2e-5
Epochs 8 3 8 3 3 8 3 8 8
LR scheduler Linear
Warmup ratio 0.1
Max sequence length 256
$\rho$ (for SAM and RW-SAM) 0.05